OpenStudy (anonymous):

1) Solve the polynomial equation 4x^4=5x^2+9 by factoring A: {3/2, i} B: {3/2, -3/2. i, -1} C:{3/2.-3/2,1,-1} 2)Simplify the expression (x-y)(x^2+xy+y^2) A: x^3-y^3 B:x^3-2xy+y^3 C:x^3+y^3 3)A family is traveling by car on a trip and they set their odometer on the car to 0 miles before they left home. After 2 hours in the car, they had traveled 80 miles. When they stopped after 10 hours of tavel, they had driven a total of 496 miles, according to the odometer reading. What is their average rate of speed for their car trip that day? A: 52mph B:60mph C:56mph Please help!!! thank you!

OpenStudy (campbell_st):

well start by rewriting the equation \[4x^4 - 5x^2 - 9 = 0\] this is an a variation on the quadratic.... so start by multiplying the coefficient of the leading term and the constant 4 * -9 = -36 find the factors of -36 that add to -5, the larger factor is negative... then you need (4x^2 + factor 1)(4x^2 + factor 2) -------------------------------- = 0 4

OpenStudy (anonymous):

I think it is A. I am not quite sure. I am trying to understand!!

OpenStudy (isaiah.feynman):

I don't think that polynomial can be solved by factoring, use the quadratic formula to solve it!

OpenStudy (campbell_st):

well the problem can be solved by factoring.... and you choice of A is it a guess of from some factoring...

OpenStudy (campbell_st):

so if you followed what I posted above... the factors of -36 and -9 and 4 so you are now looking at \[\frac{(4x^2 + 4)(4x^2 - 9)}{4}\] the 1st quadratic factor has a common factor of 4 \[\frac{4(x^2 +1)(4x^2 - 9)}{4}\] remove the common factor of 4 from the numerator and denominator which gives \[(x^2 + 1)(4x^2 - 9) = 0\] so now you are solving 2 quadratics \[x^2 + 1 = 0....and......4x^2 - 9 = 0\] solving thre above quadratics will give the solution to the problem

OpenStudy (isaiah.feynman):

I got a simpler solution!!! let \[ x^{4} = u^{2}, x^{2}=u \] we rewrite the polynomial in terms of u instead of x so it becomes \[4u^{2}-5u-9 \]. Factoring this polynomial gives us \[(4u-9)(u+1)\]. So \[u=\frac{ 9 }{ 4 }, u=-1\]. Remembering that we need x values not u values we write u as \[x^{2} \]. \[x^{2}=\frac{ 9 }{ 4 }, -1 \]. Solving for x gives us \[\frac{ 3 }{ 2 } \] and i.

OpenStudy (campbell_st):

lol..@Isaiah.Feynman funny thing ... you said it couldn't be factorised...

OpenStudy (isaiah.feynman):

I didn't look at it properly at first.